# Projective reflection groups of finite covolume

### v1

@@ 7. Minimality of the Vinberg domain
Let $\Delta=\bigcap_{s\in S}\{\alpha_{s}\leq 0\}$ be the preferred lift of $P$. By Remark 3.12 applied to the transpose of $A_{S_{f}}$, we can find a vector $\Lambda=(\lambda_{s})_{s\in S_{f}}$ with positive coordinates such that $\Lambda A_{S_{f}}<0$. In other words, if we let $\alpha=\sum_{s\in S_{f}}\lambda_{s}\alpha_{s}$, then $\alpha(v_{t})<0$ for all $t\in S_{f}$. Therefore $\alpha(v_{t})\leq 0$ for all $t\in S$ with equality if and only if $t\in S_{f}^{\perp}$. But since
$\Delta\cap\operatorname{Span}f\subset\sum_{s\in S}\mathbb{R}_{\geq 0}v_{s}\subset\{\alpha\leq 0\}$
and $\Delta\cap\operatorname{Span}f\subset\{\alpha=0\}$, it follows that
@@
i.e. $f\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S_{f}^{\perp}})$.
Suppose now that $f$ is minimal, and pick any $g\in\Sigma(f)$. By Proposition 6.3, if $S_{g}\cap S_{f}^{\perp}=\varnothing$, then $g$ is a smaller face of $P$ of negative type, contradicting the minimality of $f$. Hence the facets of $f$ are determined by the elements of $S_{f}^{\perp}$, so $(\alpha_{s},v_{s})_{s\in S_{f}^{\perp}}$ is a Coxeter polytope structure for $f\subset\operatorname{\mathbb{P}}(\operatorname{Span}f)$. Since $\Omega_{f}$ is a face of $\Omega_{P}$ which is properly convex, $\Omega_{f}$ is properly convex in $\operatorname{\mathbb{P}}(\operatorname{Span}f)$, hence $f$ is a Coxeter polytope of negative type by Proposition 4.3. Moreover, observe that if $f$ (seen as a Coxeter polytope) had a proper face $g$ of negative type, then $S_{g}\cap S_{f}^{\perp}$ would be of negative type, so $g$ would also be a negative type face of $P$ by Proposition 6.3, contradicting the minimality of $f$. Hence $f$ has no negative type faces, and is quasiperfect by Theorem 6.1. ∎
As a corollary, we obtain the following characterization of the Coxeter polytopes whose Vinberg domain is the convex hull of the proximal limit set.
Corollary 7.3.
Let $P$ be a Coxeter polytope of negative type. Then $P$ is the join of finitely many quasiperfect Coxeter polytopes of negative type if and only if $\Omega_{P}=\operatorname{int}\operatorname{Conv}(\Lambda_{P})$.
@@
If $P$ is the join of the quasiperfect Coxeter polytopes of negative type $(P_{i})_{i=1}^{k}$ corresponding to a decomposition of $V$ into the invariant direct sum $V_{1}\oplus\dots\oplus V_{k}$, then $\Gamma_{P}$ restricts to $\Gamma_{P_{i}}$ on each summand $V_{i}$. Hence $\Gamma_{P}$ is isomorphic to $\Gamma_{P_{1}}\times\dots\times\Gamma_{P_{k}}$, with $\Gamma_{P_{i}}$ acting naturally on $\Omega_{P_{i}}$ and trivially on the other factors $\Omega_{P_{j}}$ for $j\neq i$. In particular, a proximal element in $\Gamma_{P}$ is of the form $(\gamma_{1},\dots,\gamma_{k})$ where $\gamma_{i}$ is proximal for some $i$ with top eigenvalue larger than the modulus of the complex eigenvalues of the $\gamma_{j}$ for $j\neq i$, and its attractive fixed point in $\operatorname{\mathbb{P}}(V)$ is exactly the attractive fixed point of $\gamma_{i}$ in $\operatorname{\mathbb{P}}(V_{i})\subset\operatorname{\mathbb{P}}(V)$. Therefore $\Lambda_{P}=\bigcup_{i=1}^{k}\Lambda_{P_{i}}$, and since we have $\Omega_{P_{i}}=\operatorname{int}\operatorname{Conv}(\Lambda_{P_{i}})$ for each $i$ by Theorem 7.2, it follows that
$\overline{\Omega_{P}}$ $=\overline{\Omega_{P_{1}}}\mathbin{\otimes}\dots\mathbin{\otimes}\overline{\Omega_{P_{k}}}$
$=\operatorname{Conv}(\Lambda_{P_{1}})\mathbin{\otimes}\dots\mathbin{\otimes}\operatorname{Conv}(\Lambda_{P_{k}})$
$=\operatorname{Conv}\left(\Lambda_{P_{1}}\cup\dots\cup\Lambda_{P_{k}}\right)$
$=\operatorname{Conv}(\Lambda_{P})$
hence $\Omega_{P}=\operatorname{int}\operatorname{Conv}(\Lambda_{P})$.
Suppose now that $\Omega_{P}=\operatorname{int}\operatorname{Conv}(\Lambda_{P})$. By Proposition 3.14, $\Lambda_{P}\subset\operatorname{\mathbb{P}}(\operatorname{Span}(v_{s})_{s\in S})$, so $\operatorname{Span}(v_{s})_{s\in S}=V$. Observe that if $P$ is irreducible, then Theorem 7.2 gives the desired result. So we may suppose that $P$ is reducible, in which case Proposition 3.13 shows that $W_{P}$ is a reducible Coxeter group since $\rho_{P}$ is reduced and dual-reduced. By [Vin71, Cor. 4] (see [Mar17, Thm. 2.12] for an easier formulation), it follows that $P$ is decomposable.
Write $P$ as the join of indecomposable Coxeter polytopes $(P_{i})_{i=1}^{k}$ associated to a decomposition of $V$ into the invariant direct sum $V_{1}\oplus\dots\oplus V_{k}$. We also write $S=S_{1}\sqcup\dots\sqcup S_{k}$ the decomposition of $S$ associated to the decomposition of $A_{P}$ into diagonal blocks $A_{P_{1}},\dots,A_{P_{k}}$. Since $P$ is of negative type and $A_{P}$ is the block diagonal matrix with blocks $A_{P_{i}}$, it follows that each $P_{i}$ is of negative type. By the same argument as above, $\Lambda_{P}=\bigcup_{i=1}^{k}\Lambda_{P_{i}}$, and since $\Omega_{P}=\operatorname{int}\operatorname{Conv}(\Lambda_{P})$, it follows that $\Omega_{P_{i}}=\operatorname{int}\operatorname{Conv}(\Lambda_{P_{i}})$ for all $i$. Therefore, by Proposition 3.14 again, $\operatorname{Span}(v_{s})_{s\in S_{i}}=V_{i}$, hence $P_{i}$ is irreducible for each $i$. Hence, Theorem 7.2 applied to each $P_{i}$ shows that they all are quasiperfect Coxeter polytopes of negative type. ∎
@@
If $P$ is a quasiperfect Coxeter polytope such that $|S|\geq 3$, it is irreducible by Proposition 4.9, so by Proposition 3.13, $P$ is reduced and dual-reduced. Hence, Proposition 4.10 ensures that $\Omega_{P}$ is the largest $\Gamma_{P}$-invariant properly convex domain in $\operatorname{\mathbb{P}}(V)$. In particular, any $\Gamma_{P}$-invariant properly convex domain lies in $\Omega_{P}$, so by Proposition 2.6, it must contain $\operatorname{int}\operatorname{Conv}(\Lambda_{P})$. Thus $\operatorname{int}\operatorname{Conv}(\Lambda_{P})$ is in fact the smallest $\Gamma_{P}$-invariant properly convex domain of $\operatorname{\mathbb{P}}(V)$, and Theorem 7.2 implies that it equals $\Omega_{P}$, so $\Omega_{P}$ is the unique $\Gamma_{P}$-invariant properly convex domain.
On the other hand, suppose that $\Omega_{P}$ is the unique $\Gamma_{P}$-invariant properly convex domain in $\operatorname{\mathbb{P}}(V)$. In particular, we must have that $\operatorname{int}\operatorname{Conv}(\Lambda_{P})=\Omega_{P}$, so Corollary 7.3 implies that $P$ is a join of quasiperfect Coxeter polytopes of negative type. However, if $P$ is decomposable, then $\Omega_{P}$ is a join, and by the discussion following paragraph 4.2, there must be at least one more $\Gamma_{P}$-invariant properly convex domain. Hence $P$ is a quasiperfect Coxeter polytope of negative type .
Suppose that $|S|\leq 2$. Since $P$ is of negative type, we must have that $|S|=2$, $\Omega_{P}$ is a segment and $d=1$. However in this case, $\Gamma_{P}$ also preserves the properly convex complement $\operatorname{\mathbb{P}}(V)\smallsetminus\Omega_{P}$, which is a contradiction. Hence $|S|\geq 3$. ∎
@@
Theorem 7.2 can be seen as a first step in proving the result for irreducible reflection groups. Indeed, by Proposition 4.5 and the fact that if $P$ is a Coxeter polytope of negative type and $W_{P}$ is affine of type $\tilde{A}_{d}$, then $\Omega_{P}$ is a simplex (see [MV00]), we have
@@
Let $P$ be a quasiperfect Coxeter polytope of negative type. Then the extreme points of $\overline{\Omega_{P}}$ are dense in $\partial\Omega_{P}$ unless $W_{P}$ is affine of type $\tilde{A}$.

### v2

@@ 7. Minimality of the Vinberg domain
Let $\Delta=\bigcap_{s\in S}\{\alpha_{s}\leq 0\}$ be the preferred lift of $P$. By Remark 3.12 applied to the transpose of $A_{S_{f}}$, we can find a row vector $\Lambda=(\lambda_{s})_{s\in S_{f}}$ with positive coordinates such that $\Lambda A_{S_{f}}<0$. In other words, if we let $\alpha=\sum_{s\in S_{f}}\lambda_{s}\alpha_{s}$, then $\alpha(v_{t})<0$ for all $t\in S_{f}$. Therefore $\alpha(v_{t})\leq 0$ for all $t\in S$ with equality if and only if $t\in S_{f}^{\perp}$. Since by assumption, $\Delta\subset\sum_{s\in S}\mathbb{R}_{\geq 0}v_{s}\subset\{\alpha\leq 0\}$, intersecting with $\{\alpha=0\}$ yields
@@
as $\Delta\cap\{\alpha=0\}=\Delta\cap\operatorname{Span}f$. In other words, $f\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S_{f}^{\perp}})$.
Suppose now that $f$ is minimal, and pick any $g\in\Sigma(f)$. By Proposition 6.3, if $S_{g}\cap S_{f}^{\perp}=\varnothing$, then $g$ is a smaller face of $P$ of negative type, contradicting the minimality of $f$. Hence the facets of $f$ are determined by the elements of $S_{f}^{\perp}$, so $(\alpha_{s},v_{s})_{s\in S_{f}^{\perp}}$ is a Coxeter polytope structure for $f\subset\operatorname{\mathbb{P}}(\operatorname{Span}f)$. Since $\Omega_{P}$ is properly convex and the Vinberg domain $\Omega_{f}\subset\operatorname{\mathbb{P}}(\operatorname{Span}f)$ of $f$ is a subset of $\overline{\Omega_{P}}$, $\Omega_{f}$ is properly convex in $\operatorname{\mathbb{P}}(\operatorname{Span}f)$, so $f$ is a Coxeter polytope of negative type by Fact 4.3. Moreover, observe that if $f$ (seen as a Coxeter polytope) had a proper face $g$ of negative type, then by definition, $S_{g}\cap S_{f}^{\perp}$ would be of negative type, so $g$ would also be a negative type face of $P$ by Proposition 6.3, contradicting the minimality of $f$. Hence $f$ has no proper negative type faces, so it is quasiperfect by Theorem 6.1. ∎
Lemma 7.4.
Let $P$ be a Coxeter polytope of negative type such that
$P\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S})$
Then $P$ is the join of its minimal negative type faces.
@@
Let $f$ be a face of $P$ maximal for inclusion among the faces that are the join of minimal negative type faces. Since $P$ is of negative type, it always has at least one minimal negative type face, which is a (trivial) join of minimal negative type faces, so $f$ is well-defined. Suppose by contradiction that $f$ is a proper face of $P$. We will show that we can find a minimal negative type face $g$ of $P$ such that the join of $f$ and $g$ is a face of $P$, contradicting the maximality of $f$.
Write $f$ as the join of minimal negative type faces $(f_{i})_{i=1}^{k}$ of $P$. We consider
$Q:=\bigcap_{s\in S_{f}\cup S_{f}^{\perp}}\operatorname{\mathbb{P}}\{\alpha_{s}\leq 0\}$
Observe that $Q\supset P$ is non-empty and is properly convex as $\bigcap_{s\in S_{f}\cup S_{f}^{\perp}}\{\alpha_{s}=0\}=\varnothing$ by applying Remark 6.2 to $f$. Indeed, the facets of $f$ are determined by the elements of $S_{f}^{\perp}$ since $f$ is the join of the $f_{i}$ and for each $i$, the facets of $f_{i}$ are given by $S_{f_{i}}^{\perp}\subset S_{f}^{\perp}$ by Lemma 7.3. Moreover, we can endow $Q$ with a Coxeter polytope structure with the data $(\alpha_{s},v_{s})_{s\in S_{f}\cup S_{f}^{\perp}}$.
By Lemma 5.4 applied to the polytope $Q$ with $T_{1}=S_{f}^{\perp}$ and $T_{2}=S_{f}$, $S_{f}^{\perp}$ defines a face $f_{Q}^{\perp}$ of $Q$. Since $\operatorname{Span}f=\bigcap_{s\in S_{f}}\ker\alpha_{s}$ and the facets of $f$ are determined by the elements of $S_{f}^{\perp}$, $f$ is also a face of $Q$. It follows that $Q$ is the join of $f$ and $f_{Q}^{\perp}$.
Since $S_{f}$ is of negative type, we can find a vector $\Lambda:=(\lambda_{s})_{s\in S_{f}}>0$ such that $A_{S_{f}}\Lambda<0$ by Remark 3.12. If we let $x=\sum_{s\in S_{f}}\lambda_{s}v_{s}$, we then have $\alpha_{t}(x)\leq 0$ for all $t\in S$ with equality if and only if $t\in S_{f}^{\perp}$. Hence $S_{f}^{\perp}$ also defines a face $f^{\perp}:=P\cap f_{Q}^{\perp}$ of $P$ such that $S_{f^{\perp}}=S_{f}^{\perp}$ by Definition 5.3.(2).
We claim that $S_{f}^{\perp}=\bigsqcup_{i=1}^{k}S_{f_{i}}^{\perp}$. Observe that the $S_{f_{i}}^{\perp}$ are pairwise disjoint as $S_{f_{i}}^{\perp}\subset S_{f_{j}}$ whenever $j\neq i$, since $f$ is the join of the $f_{i}$ whose facets are given by $S_{f_{i}}^{\perp}$. Hence the right-hand side is indeed a disjoint union, and is clearly included in $S_{f}^{\perp}$. Moreover, the hyperplane spanned by an element of $S_{f}^{\perp}$ must define a facet of $f$ since it already contains $\operatorname{\mathbb{P}}(\operatorname{Span}f_{Q}^{\perp})$ and $Q$ is the join of $f$ and $f_{Q}^{\perp}$. Thus the reverse inclusion follows from the fact that the facets of $f$ are exactly given by the $S_{f_{i}}^{\perp}$.
Now, since $f$ is the join of finitely many Coxeter polytopes of negative type whose Coxeter polytope structure is the restriction of the Coxeter polytope structure of $P$, $f$ is also a Coxeter polytope of negative type with the Coxeter structure given by the restriction of that of $P$ to $S_{f}^{\perp}$. Therefore, $A_{S_{f}^{\perp}}$ is of negative type, so $f^{\perp}$ is a negative type face of $P$, and we can find a minimal negative type face $g$ of $P$ contained in $f^{\perp}$.
Note that $S_{f}^{\perp\perp}=S_{f}$ in $S$. This is because $s\in S_{f}^{\perp\perp}$ if and only if $\alpha_{s}(v_{t})=0$ for all $t\in S_{f}^{\perp}$, if and only if $f\subset\operatorname{\mathbb{P}}(\ker\alpha_{s})$ as $\operatorname{Span}f=\operatorname{Span}(v_{s})_{s\in S_{f}^{\perp}}$ by Lemma 7.3, if and only if $s\in S_{f}$. It follows that $f^{\perp}=f_{Q}^{\perp}$, as $\operatorname{Span}f^{\perp}=\operatorname{Span}f_{Q}^{\perp}$ and $f^{\perp}$ is determined by the elements of $S_{f^{\perp}}^{\perp}=S_{f}^{\perp\perp}=S_{f}$ by Lemma 7.3.
Therefore $Q$ is the join of $f$ and $f^{\perp}$, so the join $f\mathbin{\otimes}g$ of $f$ and $g$ is a face of $Q$ contained in $P$, as both $f$ and $g$ are. Thus $f\mathbin{\otimes}g$ is a face of $P$ since $P\subset Q$, larger than $f$, and that is the join of minimal negative type faces of $P$, which is a contradiction. Hence $f=P$, so $P$ is the join of its minimal negative type faces. ∎
@@
Suppose first that $\Omega_{P}=\operatorname{int}\operatorname{Conv}(\Lambda_{P})$. By Fact 3.14, $\Lambda_{P}\subset\operatorname{\mathbb{P}}(V_{v})$, so $V_{v}=V$ and $\rho$ is reduced and dual-reduced. If $P$ is irreducible, $\Gamma_{P}$ is either large or affine of type $\tilde{A}_{d}$ by Fact 4.5. If $\Gamma_{P}$ is affine of type $\tilde{A}_{d}$, then it is perfect, so $P\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S})$ by Lemma 7.2. If $\Gamma_{P}$ is large, Fact 4.12 and Remark 4.13 imply that $P\subset P\cap\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S})$, so $P\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S})$.
If $P$ is reducible, then $P$ is decomposable by [Vin71, Cor. 4] (see [Mar17, Thm. 2.12] for an easier formulation) since it is reduced and dual-reduced. Write $P$ as the join of irreducible Coxeter polytopes $(P_{i})_{i=1}^{k}$ associated to the $\Gamma_{P}$-invariant decomposition $V=\bigoplus_{i=1}^{k}V_{i}$. Since $P$ is of negative type, so are the $P_{i}$, and by Lemma 4.7, $\Gamma_{P}$ is proximal and $\Lambda_{P}=\bigcup_{i=1}^{k}\Lambda_{P_{i}}$. It follows that $\Omega_{P_{i}}=\operatorname{int}\operatorname{Conv}(\Lambda_{P_{i}})$ for all $i$ since $\Omega_{P}=\operatorname{int}\operatorname{Conv}(\Lambda_{P})$, so by the irreducible case above, we have $P_{i}\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S_{i}})$ for all $i$. Since $S=\bigsqcup_{i=1}^{k}S_{i}$, this implies $P\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S})$. Hence (i) implies (ii).
For (ii) implies (iii), suppose that $P\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S})$. By Lemmas 7.3 and 7.4, $P$ is the join of its minimal negative type faces, which are quasiperfect Coxeter polytopes of negative type.
Lastly, for (iii) implies (i), write $P$ as a join of the quasiperfect Coxeter polytopes of negative type $(P_{i})_{i=1}^{k}$. By [Mar17, Thm. 8.2] and its proof, we have $\Omega_{P_{i}}=\operatorname{int}\operatorname{Conv}(\Lambda_{P_{i}})$ for all $i$. Since $\Omega_{P}$ is a join of the $\Omega_{P_{i}}$, and $\Lambda_{P}$ the union of the $\Lambda_{P_{i}}$ by Lemma 4.7, it follows that $\Omega_{P}=\operatorname{int}\operatorname{Conv}(\Lambda_{P})$. ∎
The proof of Theorem 1.3 follows.
Proof of Theorem 1.3.
If $P$ is a quasiperfect Coxeter polytope such that $\dim P\geq 2$, we have $|S|\geq 3$. Since it is also irreducible by Fact 4.10, and hence reduced and dual-reduced by Fact 3.13, Fact 4.11 ensures that $\Omega_{P}$ is the largest $\Gamma_{P}$-invariant properly convex domain in $\operatorname{\mathbb{P}}(V)$. In particular, any $\Gamma_{P}$-invariant properly convex domain lies in $\Omega_{P}$, so by Fact 2.6, it must contain $\operatorname{int}\operatorname{Conv}(\Lambda_{P})$. Thus $\operatorname{int}\operatorname{Conv}(\Lambda_{P})$ is in fact the smallest $\Gamma_{P}$-invariant properly convex domain of $\operatorname{\mathbb{P}}(V)$. By Theorem 7.1, it equals $\Omega_{P}$, so $\Omega_{P}$ is the unique $\Gamma_{P}$-invariant properly convex domain.
On the other hand, suppose that $\Omega_{P}$ is the unique $\Gamma_{P}$-invariant properly convex domain in $\operatorname{\mathbb{P}}(V)$. In particular, we must have $\operatorname{int}\operatorname{Conv}(\Lambda_{P})=\Omega_{P}$, so Theorem 7.1 implies that $P$ is a join of quasiperfect Coxeter polytopes of negative type. However, if $P$ is a non-trivial join, then so is $\Omega_{P}$, and by the discussion following paragraph 4.2, there must be at least one more $\Gamma_{P}$-invariant properly convex domain. Hence $P$ is a quasiperfect Coxeter polytope of negative type.
Suppose that $\dim P\leq 1$. Since $P$ is of negative type, we must have that $\dim P=1$, so $\Omega_{P}$ is a segment in $\mathbb{RP}^{1}$. However in this case, $\Gamma_{P}$ also preserves the properly convex complement $\mathbb{RP}^{1}\smallsetminus\Omega_{P}$, which is a contradiction. Hence $\dim P\geq 2$. ∎
@@
Theorem 7.1 can be seen as a first step in proving the result for irreducible reflection groups. Indeed, by Fact 4.5 and the fact that if $P$ is a Coxeter polytope of negative type and $W_{P}$ is affine of type $\tilde{A}_{d}$, then $\Omega_{P}$ is a simplex (see [MV00]), we have
@@
Let $P$ be a quasiperfect Coxeter polytope of negative type. Then the extreme points of $\overline{\Omega_{P}}$ are dense in $\partial\Omega_{P}$ unless $W_{P}$ is affine of type $\tilde{A}_{d}$.